# An Empirical Study of Open Edge Computing Platforms: Ecosystem, Usage, and Security Risks

### v1

@@ I Introduction
However, these OECPs are indeed black boxes since little is publicly available regarding their technical mechanisms or real-world usage. Furthermore, it is intuitive to wonder what practical security/privacy implications these OECPs may render, especially considering the claimed extensive adoption by online services. For instance, our preliminary investigation reveals that none of these OECPs enforce strict vetting for edge node operators, and any third party, trusted or not, can turn their devices of heterogeneous types into edge nodes, which can introduce byzantine or even malicious edge nodes into the ecosystem of OECPs. It also breaks the traditional threat model for edge computing wherein edge nodes are fully controlled by the ECP and are usually considered as trusted. To clear away the dense mist surrounding OECPs, we conduct an extensive security study on two representative ones, namely, TipTime [10] and OneThingCloud [9]. As elaborated in §II, these two stand out in a comparative analysis of all identified OECPs, e.g., OneThingCloud advertises the largest volume of edge nodes among all OECPs. Our study would not be possible without the design and implementation of a set of automatic or semi-automatic tools. Particularly, to capture edge activities and edge traffic, a deployment framework is designed to automatically deploy edge nodes across OECPs at different locations and capture their edge traffic. Then, given edge traffic that is of multiple TBs, an efficient edge traffic analyzer is built up to concurrently process edge traffic flows, group them into pre-defined traffic categories, as well as correlating them with various real-world parties. Also, given a considerable understanding of the two OECPs, we have systematically reasoned about their security and privacy risks. To further demonstrate the identified potential risks without incurring ethical concerns, a security testbed is carefully designed to conduct attack experiments while avoiding interfering with real-world edge computing activities. As the results, our study has revealed a set of previously unknown findings for such OECPs with regards to technical mechanisms, the landscape of edge nodes, real-world usage and adoption, as well as security and privacy risks. And our key findings are highlighted as below.
@@
First of all, Both OECPs share many similarities in their technical mechanism and service model. Briefly, both enable the co-deployment of different edge computing tasks on a single edge node and each edge computing task runs like an independent kingdom inside the edge node. Also, when it comes to edge node recruitment, both support a variety of computing devices including mobile phones, desktop/laptop computers, computing servers, WIFI hotspots and routers, or even network attached storage devices. And the support of so many heterogeneous device types has likely contributed to the large scale of edge nodes. Also, the ecosystem of OECPs involve a complicated interactions among three types of participants: edge node operators, OECP providers, and tenants of edge computing tasks. Among these participants, edge node operators transform their devices into edge nodes and contribute computing/bandwidth resources for revenue. Then, ECP providers are responsible for coordinating edge nodes and distributing edge computing tasks across edge nodes, while edge computing tenants subscribe to OECPs so as to deploy edge computing tasks.
@@ III-B The Edge Traffic Analyzer
Leveraging a combination of these tricks and techniques, we have gained with high confidence a good understanding of both OECPs with regards to their technical mechanisms (i.e., network protocols), edge computing tasks, and the signatures to group traffic flows by their purposes, edge computing modules, and remote parties. Particularly, all edge computing tasks observed in our edge nodes turn out to be content delivery tasks that involve various CDN services and content providers. Also, a typical edge node tend to concurrently host multiple content delivery tasks that belong to different parties. More details will be presented in §IV.
@@ IV-C Edge-Assisted Content Delivery
CDNs and content providers. As aforementioned, all edge computing tasks observed in our study are content delivery tasks which involve the collaboration between CDN services and the open edge computing platforms. In total, we have observed the tasks of 6 different CDN services, among which, 2 are observed on the TipTime platform and 4 are on the OneThingCloud. Also, through analyzing the traffic flows of these content delivery tasks, we have identified 16 upstream content providers that subscribe to one or more of these 6 CDN services and have their content payloads delivered through edge nodes of the two OECPs. The full list of CDN services is listed in Table V along with the respective upstream content providers. Also, as aforementioned, Xingyu CDN is under direct control of the same company as OneThingCloud, while the other CDN services are independent from the two OECPs.
@@ V-C Low Threat Reputation of Edge Node IPs
In summary, as revealed by both threat platforms, a large portion of edge node IPs are concurrently involved into diverse malicious activities, which can likely render a low threat reputation and thus disrupt the legitimate edge computing tasks. For example, the CDN domain cdn.thunderstore.io was blocked because a malicious user had uploaded a crypto miner to this site [32].
@@ V-E Privacy Risks Against End Users
The information leakage to untrustful edge nodes. We also observe that end users of edge-assisted content delivery are subject to information leakage to edge nodes. The first type is the IP leakage, a typical edge node is used to deliver content on a daily base to thousands of end-user IPs that can be as far as thousands of kilometers. Besides, the daily content volume delivered by a typical edge node could exceed 15GB, as illustrated in Appendix B-D. Also, even if a content payload is encrypted during delivery, a typical edge node can still learn its plaintext. Therefore, such leakage of end-user IPs along with the knowledge of the content under delivery can enable a potential attacker to link an end-user IP address with the content payloads delivered to it, which can provide a side channel for miscreants for targeted censorship and privacy infringement. For instance, a miscreant can run edge nodes and monitor what videos the nearby residents are watching. Besides, we have also observed the leakage of the device type of end-users, which however, is limited to a subset of content delivery modules. When requesting content, some end-user devices will specify the device type by following the protocol requirements of the respective CDN module, which, along with the leaked IP address, can help an attacker locate the end user more precisely and thus incur a non-negligible privacy concern. Among content delivery flows, 461,446 are subject to the leakage of the device type and have revealed 4,367 different device types.
@@ VII Discussion
Data and code release. We have released the source code for all the tools developed to understand OECPs  [^3] https://chasesecurity.github.io/Open_Edge_Computing_Platforms/. Regarding the dataset in this paper, to protect privacy in the plaintext traffic and domains, it will be available upon request and necessary background vetting.

### v2

# Dissecting Open Edge Computing Platforms: Ecosystem, Usage, and Security Risks

@@ I Introduction
However, these OECPs are black boxes since little is publicly available regarding their technical mechanisms or real-world usage. Furthermore, it is intuitive to wonder what practical security/privacy implications these OECPs may render, especially considering the claimed extensive adoption by online services. For instance, our preliminary investigation reveals that none of these OECPs enforce strict vetting for edge node operators, and any third party, can turn their devices into edge nodes, which can introduce byzantine or even malicious edge nodes. It also breaks the traditional threat model for edge computing wherein edge nodes are fully controlled by the ECP and are usually considered as trusted. To clear away the dense mist surrounding OECPs, we conduct an extensive security study on two representative ones, namely, TipTime [10] and OneThingCloud [9]. As elaborated in §II, these two stand out in a comparative analysis of all identified OECPs. Our study would not be possible without the design and implementation of a set of automatic or semi-automatic tools. Particularly, to capture edge activities and edge traffic, a deployment framework is designed to deploy edge nodes across OECPs at different locations and capture their edge traffic. Then, given edge traffic of multiple TBs, an efficient edge traffic analyzer is built up to concurrently process edge traffic flows, group them into pre-defined traffic categories, as well as correlating them with various real-world parties. Also, a security testbed is carefully designed to empirically evaluate security risks of OECPs while avoiding interfering with real-world edge computing activities. Applying these tools leads to new understandings of OECPs with regards to their technical mechanisms, the landscape of edge nodes, real-world usage, as well as security and privacy risks. Our key findings are highlighted as below.
@@
First of all, Both OECPs share many properties in their technical mechanism and service model. Briefly, both enable the co-deployment of different edge computing tasks on a single edge node and each edge computing task runs like an independent kingdom inside the edge node. Also, when it comes to edge node recruitment, both support a variety of computing devices including mobile phones, desktop/laptop computers, computing servers, WIFI hotspots and routers, or even network attached storage devices. And the support of so many heterogeneous device types has likely contributed to the large scale of edge nodes. Also, the ecosystem of OECPs involve a complicated interactions among three types of participants: edge node operators, OECP providers, and tenants of edge computing tasks. Among these participants, edge node operators transform their devices into edge nodes and contribute computing/bandwidth resources for revenue. Then, ECP providers are responsible for coordinating edge nodes and distributing edge computing tasks across edge nodes, while edge computing tenants subscribe to OECPs so as to deploy edge computing tasks.
@@ III-B The Edge Traffic Analyzer
Leveraging a combination of these tricks and techniques, we have gained with high confidence a good understanding of both OECPs with regards to their technical mechanisms (i.e., network protocols), edge computing tasks, and the signatures to group traffic flows. Particularly, all edge computing tasks observed in our edge nodes turn out to be content delivery tasks that involve various CDN services and content providers. Also, a typical edge node tend to concurrently host multiple content delivery tasks that belong to different parties. More details will be presented in §IV.
@@ IV-C Edge-Assisted Content Delivery
CDNs and content providers. As aforementioned, all edge computing tasks observed in our study are content delivery tasks which involve the collaboration between CDN services and the open edge computing platforms. In total, we have observed the tasks of 6 different CDN services, among which, 2 are observed on the TipTime platform and 4 are on the OneThingCloud. Also, through analyzing the traffic flows of these content delivery tasks, we have identified 16 upstream content providers that subscribe to one or more of these 6 CDN services and have their content payloads delivered through edge nodes of the two OECPs. The full list of CDN services is listed in Table V along with the respective upstream content providers. Also, all CDN services except for Xingyu CDN are independent from the two OECPs.
@@ V-C Low Threat Reputation of Edge Node IPs
In summary, a large portion of edge node IPs are concurrently involved into diverse malicious activities, which can likely render a low threat reputation and thus disrupt the legitimate edge computing tasks. For example, the CDN domain cdn.thunderstore.io was blocked because a malicious user had uploaded a crypto miner to this site [32].
@@ V-E Privacy Risks Against End Users
The information leakage to untrustworthy edge nodes. We also observe that end users of edge-assisted content delivery are subject to information leakage to edge nodes. This is because a typical edge node knows not only the end-user IP addresses but also the plaintext content payloads under delivery. As an edge node in OECP can be malicious, such leakage can enable a potential attacker to link an end-user IP address with the content payloads delivered to it, which can provide a side channel for miscreants for targeted censorship and privacy infringement. For instance, a miscreant can run edge nodes and monitor what videos the nearby residents are watching. Furthermore, as learned from quantitative analysis, a typical edge node tend to deliver content to thousands of end-user IPs on a daily basis and these end-user IPs can be as far as thousands of kilometers from the edge node. What is also observed that the daily content delivery volume of a typical edge node could exceed 15GB. These quantitative results further highlight the concerning and practical impact of this information leakage. Besides, we have also observed the leakage of the device type of end-users, which however, is limited to a subset of content delivery modules. When requesting content, some end-user devices will specify the device type by following the protocol requirements of the respective CDN module, which, along with the leaked IP address, can help an attacker locate the end user more precisely and thus incur a non-negligible privacy concern. Among content delivery flows, 461,446 are subject to the leakage of the device type and have revealed 4,367 different device types.
@@ VII Discussion
Data and code release. We have released the source code for all the tools developed to capture and understand edge traffic  [^3] https://chasesecurity.github.io/Open_Edge_Computing_Platforms/. Regarding the dataset in this paper, to protect privacy in the plaintext traffic and domains, it will be available upon request and necessary background vetting.